# Analytical Emulator for the Baryon EoM inside the Fuzzy Dark Matter Soliton from Machine Learning

### v1

@@ IV FDM soliton formation with an analytical baryon EoM
We find both of $|\delta\rho(R,z)|$ and $|\delta\Phi(R,z)|$ are $\lesssim 0.04$, which is consistent with the accuracy of our analytical emulator for baryon EoM $\%4$, as shown in Fig. 3.
@@ V Summary and discussion
In this paper, we first solve the SP system with a fixed empirical baryon density profile background in cylindrical coordinates by an iteration of $\lambda$. For $m=10^{-22}\rm{eV}/c^{2}$, $\lambda\approx 6\times 10^{-7}$ rescales the dimensionless $\tilde{\rho}_{\rm b}$ and $\tilde{M}$ to the baryon density profile and the FDM soliton mass of the Milky Way simultaneously. Therefore, $\lambda\approx 6\times 10^{-7}$ was used to map the scaled dimensionless variables to the corresponding physical dimensional ones and vice versa. Then we make the obtained stationary solutions mock data and construct an analytical emulator (with $\%{\rm Acc}<\%4$) for the baryon EoM inside the FDM soliton by GA, using a machine learning technique. Finally, we complement a model validation by comparison between the stationary solutions of an enlarged SP system with the emulator $\tilde{\rho}_{\rm b,EoM}$ as the third equation under certain conditions and ones of the original SP system. We find the fractional errors of two sets of solutions are $\lesssim 0.04$ and are consistent with the accuracy of emulator for baryon EoM $\%4$.
What about its validation for the nonstationary evolutions of FDM solitons, especially those evolutions totally departing from the stationary solutions, such as FDM soliton collisions. Since the mock data used in this paper is extracted from the stationary solutions of the SP system with a fixed empirical baryon density profile background, our analytical emulator for the baryon EoM inside the FDM soliton should also be valid for the perturbation, the tidal disruption or deformation of FDM solitons conservatively. However, for some dramatic evolutions, our emulator probably does not perform correctly.

### v2

@@ IV FDM soliton formation with the baryon AE
We find both of $|\delta\rho(R,z)|$ and $|\delta\Phi(R,z)|$ are $\lesssim 0.04$, which is consistent with the accuracy of our baryon AE $4\%$, as shown in Fig. 3.
@@ V Summary and discussion
In this paper, we first solve the SP system with a fixed empirical baryon density profile background in cylindrical coordinates by an iteration of $\lambda$. For $m=10^{-22}\rm{eV}/c^{2}$, $\lambda\approx 6\times 10^{-7}$ rescales the dimensionless $\tilde{\rho}_{\rm b}$ and $\tilde{M}$ to the baryon density profile and the FDM soliton mass of the Milky Way simultaneously. Therefore, $\lambda\approx 6\times 10^{-7}$ was used to map the scaled dimensionless variables to the corresponding physical dimensional ones and vice versa. Then we make the obtained stationary solutions mock data and construct an analytical emulator (with $\%{\rm Acc}<4\%$) for the baryon density distribution inside the FDM soliton by GA, using a machine learning technique. Finally, we complement a model validation by comparison between the stationary solutions of an enlarged SP system with the emulator $\tilde{\rho}_{\rm b,AE}$ as the third equation under certain conditions and ones of the original SP system. We find the fractional errors of two sets of solutions are $\lesssim 0.04$ and are consistent with the accuracy of baryon AE $4\%$.
What about its validation for the nonstationary evolutions of FDM solitons, especially those evolutions totally departing from the stationary solutions, such as FDM soliton collisions. Since the mock data used in this paper is extracted from the stationary solutions of the SP system with a fixed empirical baryon density profile background, our baryon AE inside the FDM soliton should also be valid and can serve as a kind of baryon EoM for the perturbation, the tidal disruption or deformation of FDM solitons conservatively. However, for some dramatic evolutions, our baryon AE probably does not perform correctly.